# CORTEX-AVD: CORner Case Testing & EXploration for Autonomous Vehicles Development

### v1

@@ V Discussion
Despite growing interest in heuristic-based CC generation, standardized benchmarks remain scarce. For that reason, an effort was made to organize the literature using a set of common metrics (Number of Test Cases, Time, Number of Collisions (NC), Number of Valid Scenarios), used for GA efficiency evaluation (Table IV). However, no metrics related to near-collisions were listed, as none of these studies explicitly collected this information, making comparisons impossible. Variations across studies often reflect differing focuses, such as topology generation [65, 63, 57] or risk-related metrics like time [58, 58], distances [63, 57, 68, 59], and collision rates [65, 68, 59]. While recent studies underscore GA’s value for CC generation, most experiments involve only a few thousand simulations and focus on limited metrics [62, 65, 63, 66, 56, 57, 68, 58, 59, 60]. RL results are also highly dependent on GA modeling and simulation frameworks, making cross-study comparisons challenging.
@@
Although the study effectively generated CC data, some areas could be potentially improved. Scenic’s rejection sampling introduced variability in generating feasible scenarios, especially when restrictive parameter ranges were applied. Slightly relaxing these ranges could increase the number of high-risk simulations, enhancing the overall effectiveness of the approach. GA’s performance depended heavily on initial conditions, stopping criteria, and objective functions, which can be potentially fine tuned [62, 65, 57, 68, 58, 59, 61]. Additionally, testing focused on a single map in Carla, limiting exploration of diverse environments and broader risk factors such as overtaking, variable weather conditions, and obstacles, which affects the generalizability of the proposed approach.
VI Concluding Remarks
@@ Acknowledgments
This work was supported by Itaú Unibanco S.A., through the Itaú Scholarships Program (PBI), linked to the Data Science Center at the Polytechnic School of the University of São Paulo.

### v3

@@ VI Discussion
Despite growing interest in heuristic-based CC generation, standardized benchmarks remain scarce. For that reason, an effort was made to organize the literature using a set of common metrics (Number of Test Cases, Time, Number of Collisions (NC), Number of Valid Scenarios), used for GA efficiency evaluation (Table IV). However, no metrics related to near-collisions were listed, as none of these studies explicitly collected this information, making comparisons impossible. Variations across studies often reflect differing focuses, such as topology generation [66, 64, 58] or risk-related metrics like time [59, 62, 63], distances [64, 58, 69, 60], and collision rates [66, 69, 60]. While recent studies underscore GA’s value for CC generation, most experiments involve only a few thousand simulations and focus on limited metrics [63, 66, 64, 67, 57, 58, 69, 59, 60, 61]. RL results are also highly dependent on GA modeling and simulation frameworks, making cross-study comparisons challenging.
@@
Although the framework effectively generated CC data, some areas could be potentially improved. Scenic’s rejection sampling introduced variability in generating feasible scenarios, especially when restrictive parameter ranges were applied. Slightly relaxing these ranges could increase the number of high-risk simulations, enhancing the overall effectiveness of the approach. GA’s performance depended heavily on initial conditions, stopping criteria, and objective functions, which can be potentially fine tuned [63, 66, 58, 69, 59, 60, 62]. Additionally, testing focused on a single map in Carla, limiting exploration of diverse environments and broader risk factors such as overtaking, variable weather conditions, and obstacles, which affects the generalizability of the proposed approach.
VII Concluding Remarks
@@ Acknowledgments
The Article Processing Charge (APC) for the publication of this research was funded by the Coordenaçãao de Aperfeiçoamento de Pessoal de Nível Superior-Brasil (CAPES). Gabriel Kenji Godoy Shimanuki was supported by Itaú Unibanco S.A., through the Itaú Scholarships Program (PBI), linked to the Data Science Center at the Polytechnic School of the University of São Paulo.